# Exploring Representations and Interventions in Time Series Foundation Models

### v3

@@ Acknowledgment
This work was partially supported by NSF (awards 2406231 and 2427948), NIH (awards R01NS124642, R01DK131586, and R01NR013912), and the US Army (W911NF-20-D0002).
Impact Statement
While our pruned TSFMs demonstrate promising performance, it is crucial to exercise caution when using them, especially in high-stakes applications such as healthcare. Before deploying these models for critical decision-making, we strongly recommend fine-tuning and evaluating them on task-specific, in-domain datasets using relevant metrics. The ability to steer model predictions offers numerous benefits but also raises concerns regarding potential biases. We urge users to exercise caution when utilizing our proposed steering strategies and to carefully consider the potential implications of manipulating model outputs.
Reproducibility Statement
To ensure reproduciblity, we have made our code anonymously accessible through https://anonymous.4open.science/r/tsfm-interventions-3452/. All time series foundation models used for our analysis are publicly available and open-source: Chronos (https://github.com/amazon-science/chronos-forecasting), MOMENT (https://github.com/moment-timeseries-foundation-model/moment), and MOIRAI (https://github.com/SalesforceAIResearch/uni2ts). All models were trained and evaluated on a computing cluster consisting of 128 AMD EPYC 7502 CPUs, 503 GB of RAM, and 8 NVIDIA RTX A6000 GPUs each with 49 GiB RAM. Synthetic datasets used in our study and the pruned time series foundation models will be released publicly upon paper acceptance.
@@ Appendix C Finding and Pruning Redundant Blocks in TSFMs
Our pruning algorithm is summarized in Algorithm 1. Our pruning method involves retaining the first and last layers of each block while zeroing out the weights of the intermediate layers. This approach preserves the structural integrity of the block while leveraging the skip connections within transformer blocks to ensure that signals and gradients continue to flow through the network. In the case of MOMENT-Large’s Block 3, composed of layers 19–23, this means that layers 19 and 23 are retained, while the weights of layers 20, 21, and 22 are zeroed out. We have added a dedicated section on pruning in Appendix C to further clarify this process.
@@ E.3 Moirai
Moirai is a time-series foundation model built with an encoder-only Transformer architecture designed for universal forecasting. The model handles varying temporal resolutions with multiple patch size projection layers and uses any-variate attention for multivariate time series. A mixture distribution is employed to model probabilistic forecasts.
  • Architecture: Encoder-only Transformer with any-variate attention for multivariate time-series forecasting.
@@
Appendix F Additional Results
F.1 Steering
@@
Refer to caption Refer to caption
(i) Periodicity, $\beta=0$ (ii) Some trend, more periodicity, $\beta=0.25$
Refer to caption Refer to caption
(iii) Trend and periodicity, $\beta=0.5$ (iv) More trend, some periodicity, $\beta=0.75$
(v) Trend, $\beta=1.0$
Figure 12: Compositional Steering for Forecasting using the MOMENT model. The parameter $\beta$ influences compositional steering, where both sinusoidal and increasing trend concepts are observed in the steered output when $\beta$=0.5, showcasing how the steering technique can interpolate between different concept combinations in the latent space (ii). Steered output changes from sinusodial concepts to increasing trend concepts by varying $\beta\in\{0.0,0.25,0.5,0.75,1.0\}$.
Refer to caption Refer to caption
(i) Periodicity, $\beta=0$ (ii) Some trend, more periodicity, $\beta=0.25$
Refer to caption Refer to caption
(iii) Trend and periodicity, $\beta=0.5$ (iv) More trend, some periodicity, $\beta=0.75$
(v) Trend, $\beta=1.0$
Figure 13: Compositional Steering for Forecasting using the Chronos model. The parameter $\beta$ influences compositional steering, where both sinusoidal and increasing trend concepts are observed in the steered output when $\beta$=0.5, showcasing how the steering technique can interpolate between different concept combinations in the latent space (ii). Steered output changes from sinusoidal concepts to increasing trend concepts by varying $\beta\in\{0.0,0.25,0.5,0.75,1.0\}$.
@@
Figure 14: Visualization of Concept Steering on ECG Data. The figure shows PCA projections of the activation space, where arrows indicate the direction of steering. Blue and red points represent normal (class 0) and abnormal (class 1) samples respectively, with green points showing the original samples and orange points showing their steered versions.
F.2 Pruning
@@
Refer to caption Refer to caption Refer to caption Refer to caption
(i) Pattern (ii) Trend (iii) Periodicity (iv) Amplitude
Figure 18: Progression of linear separability of concepts at the patch level (y-axis) across different layers (x-axis). Linear separability is measured using the Linear Discriminant Ratio (LDR), derived from model embedding statistics for each predicted class: (i) constant vs. sinusoidal patterns, (ii) increasing vs. decreasing trends, (iii) high vs. low periodicity and (iv) high vs. low amplitude. Color gradient indicates separability, with dark blue representing low LDR and yellow indicating high LDR. Certain concepts captured by MOMENT-Large exhibit linear separability, but this separability is not uniform across layers; instead, it emerges at specific points in the model.

### v4

@@ Acknowledgment
Funding  This work was partially supported by NSF (awards 2406231 and 2427948), NIH (awards R01NS124642, R01DK131586, and R01NR013912), and the US Army (W911NF-20-D0002). Part of this work was conducted during the Robotics Institute Summer Scholars (RISS) program at Carnegie Mellon University. We gratefully thank Rachel Burcin, Dr. John Dolan, and the entire RISS team for their coordination, mentorship, and support throughout the program.
Datasets and Tooling  We are grateful to the creators and maintainers of the UCR Time-Series Archives (classification & anomaly) (Dau et al., 2018; Keogh et al., 2021), the Monash Forecasting Archive team (Godahewa et al., 2021), the Informer long-horizon data providers (ETT, Electricity, Exchange Rate, Weather, Traffic, ILI) (Zhou et al., 2021), the custodians of TSB-UAD (Paparrizos et al., 2022), and every other public dataset we relied on. We would also like to thank the open-source community– authors of the libraries, frameworks, and tooling that make modern machine learning research possible.
Discussions  We also thank the anonymous reviewers and meta-reviewer for their constructive feedback and suggestions, which helped improve the quality and clarity of this work. We are grateful to Carnegie Mellon University, the Robotics Institute, and the Auton Lab for providing the research environment and infrastructure that made this work possible. We also thank our fellow Auton Lab members and RISS interns (especially Jieshi Chen, Piotr Bartosiewicz, Chi-En Teh, Nicholas Gisolfi, Ignacy Stepka and Kacper Trebacz) for the inspiring discussions and support that helped shape this work.
Miscellaneous  Special thanks to Chipotle, Starbucks, and Wushiland Boba near Craig Street/CMU; K-Station BBQ at Schenley Plaza; Pigeon Bagels; and the on-campus coffee shops – La Prima Espresso Company, De Fer Coffee & Tea, and Redhawk Coffee Roasters for keeping the authors focused, caffeinated, and well-fed throughout the development of this project.
@@ Appendix C Finding and Pruning Redundant Blocks in TSFMs
Our pruning algorithm is summarized in Algorithm 1. Our pruning method involves retaining the first and last layers of each block while zeroing out the weights of the intermediate layers. This approach preserves the structural integrity of the block while leveraging the skip connections within transformer blocks to ensure that signals and gradients continue to flow through the network. In the case of MOMENT-Large’s Block 3, composed of layers 19–23, this means that layers 19 and 23 are retained, while the weights of layers 20, 21, and 22 are skipped. We have added a dedicated section on pruning in Appendix C to further clarify this process.
@@ E.3 Moirai
Moirai is a time series foundation model built with an encoder-only Transformer architecture designed for universal forecasting. The model handles varying temporal resolutions with multiple patch size projection layers and uses any-variate attention for multivariate time series. A mixture distribution is employed to model probabilistic forecasts.
  • Architecture: Encoder-only Transformer with any-variate attention for multivariate time series forecasting.
@@
Appendix F Additional Steering Results
Refer to caption Refer to caption
(i) Periodicity, $\beta=0$
Refer to caption Refer to caption
(ii) Some trend, more periodicity, $\beta=0.25$
Refer to caption Refer to caption
(iii) Trend and periodicity, $\beta=0.5$
Refer to caption Refer to caption
(iv) More trend, some periodicity, $\beta=0.75$
Refer to caption Refer to caption
(v) Trend, $\beta=1.0$
Figure 10: Visualization of the MOMENT model’s reconstruction (left) and the Chronos model’s forecasting (right) predictions, comparing concept steering in the latent space (blue) with the unsteered baseline (orange). The steering parameter $\beta$ can be used to interpolate between sinusodial concepts to increasing trend concepts by varying $\beta\in[0.0,1.0]$.
@@ Appendix F Additional Steering Results
Figure 13: Visualization of the impact of concept steering on the activation space of the ECG data model. The figure shows PCA projections of the activation space, where arrows indicate the direction of steering. Blue and red points represent normal (class 0) and abnormal (class 1) samples respectively, with green points showing the original samples and orange points showing their steered versions.
Refer to caption Refer to caption Refer to caption Refer to caption
(i) Pattern (ii) Trend (iii) Periodicity (iv) Amplitude
Figure 14: Linear separability of concepts at the patch level (y-axis) across model layers (x-axis). Yellow indicates higher separability than blue in the colorbar. Linear separability is measured using the Linear Discriminant Ratio (LDR), derived from model embedding statistics for each predicted class: (i) constant vs. sinusoidal patterns, (ii) constant vs. increasing trend patterns, (iii) sinusoidal vs. increasing trend patterns and (iv) sinusoidal vs. decreasing trend patterns. Color gradient indicates separability, with dark blue representing low LDR and yellow indicating high LDR. Certain concepts captured by MOMENT-Large exhibit linear separability, but this separability is not uniform across layers; instead, it emerges at specific points in the model.
Appendix G Additional Pruning Results